In a rectangle the ratio of length to width is 9:4.The length is 24cm. What is the width?
Answer by ewatrrr(24785)

 
Hi
ratio of length to width is 9:4, length = 24cm
......
Let x represent the constant of proportionality
9x = 24cm
x = 24/9 = 8/3 cm
.......
w = 4x = 32/3 cm 0r w = 10 2/3 cm